two complementray angles have degree measurements p and 3p-6.

Answer by checkley71(8403)

p+3p-6=180 4p=180-6 or 4p=174 or p=43.5 the 3p+6=3*43.5+6=136.5
proof 43.5+136.5=180 or 180=180